OpenStudy (anonymous):

The first term and the 3rd term of a geometric sequence are 2 and 162 respectively. (a) Find the general term Tn of the sequesnce. (b) If the common ratio is less than 0, how many +ve terms are less than 15000?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

\(\Large T_n = a_1 r^{n-1}\) so, first term = a1 = 2 3rd term, n=3 T3 = 162 \(\Large 162 = 2 r^{3-1}\) firsr find 'r' from here.

OpenStudy (anonymous):

r=9

OpenStudy (anonymous):

T(2)=18 or -18?

hartnn (hartnn):

we are done with part a, right ? r= 9 or -9 Tn = 2 *(9)^{n-1} or Tn = 2* (-9)^{n-1}

hartnn (hartnn):

sorry, i would revise my statement for part b, take r = -9. (so alternate terms will be positive) so, we will find total number of terms less than 15000 and divide it by 2 to get number of positive terms, because half of the terms will be positive, and other half will be negative.

OpenStudy (anonymous):

using calculator, n=6. But i don't have the correct answer.......

hartnn (hartnn):

ok, we can do this, lets consider r as positive = +9, and find the number of terms less than 15000 when r = -9, the only difference will be that half of those terms will be negative, so we can just divide that 'n' by 2 to get total number of positive terms when r is -9 does this make sense to you ?? 2 (9)^{n-1} <15000 get n

OpenStudy (anonymous):

but it says "if the common ratio is less than 0"

OpenStudy (anonymous):

if r=+9, n=5.06088